-4x-2y=-12
4x+8y=-24

Answer by nerdybill(7384)

Given:
-4x-2y=-12
4x+8y=-24
.
applying the "addition method" we "add" both equations together:
-4x-2y=-12
4x+8y=-24
----------
0x+6y=-36
6y=-36
y=-6
.
substitute the above into:
-4x-2y=-12
-4x-2(-6)=-12
-4x-(-12)=-12
-4x+12=-12
-4x=-24
4x=24
x=6
.
answer:(x,y) = (6,-6)